OpenStudy (anonymous):

Another formula that models growth is shown, where No is the initial size of the population, N is the final size of the population, r is the rate of growth or decay per time period, and t is the number of time periods. In this case, r = 100% or 1. Using this formula, how many rabbits would Daisy have after 3 years? N=N_o(1+r)^t a) 8,192 b) 16,384 c) 4,096 d) 2,058

OpenStudy (campbell_st):

well you need daisy's initial population of rabbits and the growth rate to find the population number in 3 years.

OpenStudy (campbell_st):

so based on the information you'll have \[N = N_{o}\times (1 + 1)^3..... or ... N = N_{o} \times 2^3\]